OpenStudy (anonymous):

Find the 20th number Of 4,8,12,

OpenStudy (lgbasallote):

hmm first ask yourself this...is that arithmetic progression? or geometric?

OpenStudy (anonymous):

Geometric

OpenStudy (anonymous):

What kind of operation ?

OpenStudy (anonymous):

Multiplication

OpenStudy (lgbasallote):

hmm try this to check.. 8/4 = 12/8 ?

OpenStudy (anonymous):

Oh then addition

OpenStudy (anonymous):

If it's addition, it's just +4 every time.

OpenStudy (anonymous):

So 80.

OpenStudy (lgbasallote):

right! so the equation for arithmetic progression is... an = a1 + (n-1)d can you tell me which of these are given..which you're looking for? an = ? a1 = ? n = ? d = ?

OpenStudy (anonymous):

@Numbers Is it crystal clear to you?